OpenStudy (anonymous):

Use complete sentences to describe the steps taken to simplify this problem. Make sure you include information about the new common denominator, the final simplified expression, and any restrictions.

OpenStudy (anonymous):

\[\frac{ 3 }{ x }-\frac{ 3 }{ x+1 }+\frac{ 3 }{ x+2 }\]

OpenStudy (anonymous):

it equals (3(x+1)(x+2)-3x(x+2)+3x(x+1))/(x(x+1)(x+2)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

just consider / as fraction it will be done

OpenStudy (anonymous):

do i need to multiply or something?

OpenStudy (anonymous):

on the nominator you should multiply then you can get the last answer

OpenStudy (anonymous):

okayy i got 6x+6, is that right?

OpenStudy (anonymous):

no try it again [3(x+1)(x+2)-3x(x+2)+3x(x+1)]/(x(x+1)(x+2)

OpenStudy (anonymous):

how about 3x^2+6x+6?

OpenStudy (anonymous):

do i multiply the bottom?

OpenStudy (anonymous):

yes you got right answer do it for bottom too

OpenStudy (anonymous):

x^3+3x+2x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

just the things you got for up and bottom

OpenStudy (anonymous):

\[\frac{ 3x^2+6x+6 }{ x^3+3x^2+2x }?\]

OpenStudy (anonymous):

good

OpenStudy (anonymous):

is that simplified?

OpenStudy (anonymous):

no its just in that way no simplize

OpenStudy (anonymous):

okayy thanks